# An interpretation of the Brownian bridge as a physics-informed prior for the Poisson equation

### v1

@@ 4.4 A note on the use of neural networks
As PINNs is a motivating application in this work, we discuss the use of deep neural networks as applied to our theorems. When working with neural networks, there are some technical issues which must be treated with care. We touch on both treating the the space $\mathcal{F}$ as a set of neural networks as well as the convergence theorems.
Recall we may choose to approximate the prior with the finite-dimensional representation as given by eq. (4.2). This representation is not immediately well-defined if we parameterize $u$ with a neural network. To summarize, for a fixed neural network structure, we cannot assume that the space of functions the network can represent is finite-dimensional. To explain this, we introduce some notation following [38, 34].
Let $\Phi=\{(A_{\ell},b_{\ell}\}_{\ell=1}^{L}$ be a set of matrix-vector tuples where $A_{\ell}\in\mathbb{R}^{N_{\ell}\times N_{\ell-1}}$ and $b_{\ell}\in\mathbb{R}^{N_{\ell}}$ for each $\ell$. The architecture of the network is given by $S=(N_{0},N_{1},\dots,N_{L})$, where $N(S)$ is the number of neurons and $L=L(S)$ is the number of layers. The collection $\Phi$ represents the possible values of the weights for a neural network with architecture $S$. Then, given an activation function $h:\mathbb{R}\to\mathbb{R}$, the neural network is given by the map $\mathrm{NN}_{h}(\Phi):\Omega\to\mathbb{R}$. We are interested in the properties of the function space induced by the network for fixed $S$ and $\sigma$. We will denote this set by $\mathcal{R}(\mathrm{NN_{h}})(S)$.
As it turns out, if we allow $\Phi$ to vary arbitrarily, then $\mathcal{R}(\mathrm{NN_{h}})(S)$ is not closed in $L^{p}(\Omega)$, $0<p<\infty$, for all activation functions commonly used in PINNs [38, Theorem 3.1]. The same is true in Sobolev spaces [34]. Then, $\mathcal{F}=\mathcal{R}(\mathrm{NN_{h}})(S)$ is not finite-dimensional and eq. (4.2) is no longer well-defined.
However, if $\Phi$ is restricted to a compact set, then $\mathcal{R}(\mathrm{NN_{h}})(S)$ is compact in $L^{p}(\Omega)$ [38, Proposition 3.5]. This compact restriction of $\Phi$ results from schemes which prevent exploding weights, a common practice. While the result is nicer, it is still not immediately applicable to the construction of our finite-dimensional approximation: there is no guarantee that a compact set of a Hilbert space will be finite-dimensional. The Hilbert cube is one example. Although, we may approximate any compact set with a finite-dimensional subspace to arbitrary accuracy.
Theorem 4.8.
@@
Therefore, one could theoretically take $\mathcal{F}$ to be a finite-dimensional space which approximates $\mathcal{R}(\mathrm{NN}_{h})(S)$ to a given tolerance, $\varepsilon$. The size of $\mathcal{F}$ on which $\mu^{(n)}$ is defined may be adjusted by tweaking $\varepsilon$, changing the bound on the weights, or changing the network structure.
In Corollary 4.2, we show that the function which solves the physics-informed optimization problem will converge to the ground truth in the large-data limit. This is if we solve the problem in the infinite-dimensional setting. Ideally we would like to derive the result for training neural networks. The solution to this problem will be a function which lives in the Sobolev space $H_{k}$. Again if we allow $\Phi$ to vary arbitrarily, then $\mathcal{R}(\mathrm{NN}_{h})(S)$ is not closed in $H_{k}$. This means that there are functions in $H_{k}$ for which the neural network must send $\|\Phi\|\to\infty$ in order to approximate. If the ground truth happens to be such a function, then the convergence theorem will not hold. Likewise, if we limit $\Phi$ to a compact set, then $\mathcal{R}(\mathrm{NN}_{h})(S)$ is compact in $H_{k}$. In this case, the neural network is only able to approximate any function to any accuracy if that function is also a neural network, so it is unlikely that convergence holds. The only case where convergence to the ground truth could hold is if we allow the architecture of the neural network to change arbitrarily so that we may rely on the universal approximation theorem [27].
@@ 5 On model-form error
In this section we perform an analysis of the hyperparameter $\beta$ towards the application of detecting model-form error. Since $\beta$ is a hyperparameter of the GP prior, it is natural to assess how $\beta$ is learned during training. We show the optimal choice of $\beta$ adjusts according to model-misspecification. We build towards the result by working with the finite-dimensional distributions.
Start by introducing a finite-dimensional representation of $L^{2}(\Omega)$. This representation induces the function space $\mathcal{F}\subset L^{2}(\Omega)$ with $\dim(\mathcal{F})=M<\infty$. We will then study what happens in the limit of infinite data. Given our training data of the form eq. (2.3), we begin by writing the problem down as a hierarchical model:
$\beta$ $\sim p(\beta)$
$\hat{u}|\beta$ $\sim\mathcal{N}(\hat{u}^{0},\beta^{-1}\Sigma_{\mathcal{F}})$
@@
where $\hat{u}^{0}$ is the projection of $u^{0}$ onto $\mathcal{F}$, and $\Sigma_{\mathcal{F}}$ is the restriction of the covariance operator given by eq. (3.4). Since we are no longer in the infinite-dimensional setting, application of Bayes’s rule in the usual sense is justified, and we can also rely on the Lebesgue integral when deriving expressions. Therefore, we derive the joint posterior
@@
In what follows, we will show the MAP estimate is unique in the limit of large data, for certain choices of $p(\beta)$. We start by deriving an expression for the gradient of the target function. Throughout, we will center the space so that the prior mean function becomes 00. We have shown this is valid as the prior mean function does not depend on $\beta$ and it also lives in $H_{k}$.
Lemma 5.1.
[…]
$\frac{\partial}{\partial\beta}\mathcal{L}(\beta)=\frac{M}{2\beta}+\frac{\partial}{\partial\beta}\log p(\beta)-\frac{1}{2}\langle\tilde{m}-\hat{u}^{0},\Sigma^{-1}_{\mathcal{F}}(\tilde{m}-\hat{u}^{0})\rangle-\frac{1}{2}\mathrm{tr}\left(\Sigma^{-1}_{\mathcal{F}}\tilde{\Sigma}_{\mathcal{F}}\right),$
where $\tilde{m}(\cdot)$ and $\tilde{\Sigma}_{\mathcal{F}}:\mathcal{F\to\mathcal{F}}$ are given by the posterior mean function eq. (2.4) and posterior covariance form eq. (2.5), respectively, and $\hat{u}^{0}$ is the projection of $u^{0}$ onto $\mathcal{F}$.
@@
[…]
$\partial_{\beta}\log\int\frac{1}{Z}p(d|\hat{u})p(\hat{u}|\beta)=\frac{\partial_{\beta}\int\frac{1}{Z}p(d|\hat{u})p(\hat{u}|\beta)}{\int\frac{1}{Z}p(d|\hat{u})p(\hat{u}|\beta)}=\frac{\partial_{\beta}\int p(d|\hat{u})p(\hat{u}|\beta)}{\int p(d|\hat{u})p(\hat{u}|\beta)}$ (5.2)
In the numerator, the only term which depends on $\beta$ is $p(\hat{u}|\beta)$, so passing the derivative through the integral, and writing the density of $p(\hat{u}|\beta)$, we get
$\partial_{\beta}p(\hat{u}|\beta)$ $=\partial_{\beta}\frac{1}{(2\pi)^{M/2}|\beta^{-1}\Sigma_{\mathcal{F}}|^{1/2}}\exp\left\{-\frac{1}{2}\langle\hat{u},\beta\Sigma_{\mathcal{F}}^{-1}\hat{u}\rangle\right\}$
$=\partial_{\beta}\frac{\beta^{M/2}}{(2\pi)^{M/2}|\Sigma_{\mathcal{F}}|^{1/2}}\exp\left\{-\frac{1}{2}\langle\hat{u},\beta\Sigma_{\mathcal{F}}^{-1}\hat{u}\rangle\right\}$
$=\frac{M\beta^{M/2-1}/2}{(2\pi)^{M/2}|\Sigma_{\mathcal{F}}|^{1/2}}\exp\left\{-\frac{1}{2}\langle\hat{u},\beta\Sigma_{\mathcal{F}}^{-1}\hat{u}\rangle\right\}$
$-\frac{\beta^{M/2}/2}{(2\pi)^{M/2}|\Sigma_{\mathcal{F}}|^{1/2}}\langle\hat{u},\Sigma^{-1}_{\mathcal{F}}\hat{u}\rangle\exp\left\{-\frac{1}{2}\langle\hat{u},\beta\Sigma_{\mathcal{F}}^{-1}\hat{u}\rangle\right\}$
$=\left(\frac{M}{2\beta}-\frac{1}{2}\langle\hat{u},\Sigma_{\mathcal{F}}^{-1}\hat{u}\rangle\right)\mathcal{N}(0,\beta^{-1}\Sigma_{\mathcal{F}}).$ (5.3)
Inserting the expression in eq. (5.3) into eq. (5.2) gives
[…]
Now, observe that when we pass the term $\frac{1}{p(d|\hat{u})p(\hat{u}|\beta)}$ inside of the integral and multiply it by $\mathcal{N}(0,\beta^{-1}\Sigma_{\mathcal{F}})$, the result is $p(\hat{u}|d,\beta)$ through application of Bayes’s rule. This is the multivariate Gaussian $\mathcal{N}(\tilde{m},\tilde{\Sigma}_{\mathcal{F}})$ where $\tilde{m}$ is given by eq. (2.4) and $\tilde{\Sigma}_{\mathcal{F}}$ is the covariance form on $\mathcal{F}$ with kernel given by eq. (2.5). So, we observe that eq. (5.4) reduces to a multivariate Gaussian integral against a quadratic form, yielding:
$\partial_{\beta}\log\int\frac{1}{Z}p(d|\hat{u})p(\hat{u}|\beta)=\frac{M}{2\beta}-\frac{1}{2}\langle\tilde{m},\tilde{\Sigma}_{\mathcal{F}}\tilde{m}\rangle-\frac{1}{2}\mathrm{tr}\left(\Sigma^{-1}_{\mathcal{F}}\tilde{\Sigma}_{\mathcal{F}}\right),$
and shifting the mean back to $\hat{u}^{0}$ completes the proof. ∎
We now identify the MAP estimate of $\beta$ under different prior choices.
@@
[…]
  1. (i)
    If $\beta$ is assigned a flat prior, then
    $\beta^{*}=\frac{M}{\langle\hat{u}^{*}-\hat{u}^{0},\Sigma_{\mathcal{F}}^{-1}(\hat{u}^{*}-\hat{u}^{0})\rangle}.$
  2. (ii)
    If $\beta$ is assigned Jeffreys prior, then
    $\beta^{*}=\frac{M-2}{\langle\hat{u}^{*}-\hat{u}^{0},\Sigma_{\mathcal{F}}^{-1}(\hat{u}^{*}-\hat{u}^{0})\rangle}.$
@@
$\frac{\partial}{\partial\beta}\mathcal{L}(\beta)=\frac{M}{2\beta}+\frac{\partial}{\partial\beta}\log p(\beta)-\frac{1}{2}\langle\tilde{m}-\hat{u}^{0},\Sigma^{-1}_{\mathcal{F}}(\tilde{m}-\hat{u}^{0})\rangle-\frac{1}{2}\mathrm{tr}\left(\Sigma^{-1}_{\mathcal{F}}\tilde{\Sigma}_{\mathcal{F}}\right).$
Now, by Theorem 4.7 we have $\tilde{m}\to u^{*}$, and by Theorem 4.7 $\tilde{k}\to 0$. Passing to the limit, the gradient becomes
$\frac{\partial}{\partial\beta}\mathcal{L}(\beta)=\frac{M}{2\beta}+\frac{\partial}{\partial\beta}\log p(\beta)-\frac{1}{2}\langle\hat{u}^{*}-\hat{u}^{0},\Sigma^{-1}_{\mathcal{F}}(\hat{u}^{*}-\hat{u}^{0})\rangle.$
@@
$\langle\hat{u}^{*}-\hat{u}^{0},\Sigma_{\mathcal{F}}^{-1}(\hat{u}^{*}-\hat{u}^{0})\rangle=\|\hat{u}^{*}-\hat{u}^{0}\|_{H_{k}}^{2},$
restricted to $\mathcal{F}$, which was derived in the proof of Proposition 3.1. That is, the optimal value of $\beta$ is inversely proportional to the RKHS distance between the prior mean function $u^{0}$ and the ground truth $u^{*}$. The same holds for $L^{2}(\Omega)$ distance by the Sobolev embedding theorem. Recall that $u^{0}$ is exactly the unique solution to the chosen physical model eq. (1.1). Hence, $\beta$ is sensitive to the distance between the true field $u^{*}$, and the one we have a priori assumed is correct $u^{0}$. As $u^{*}$ moves further away from $u^{0}$, the optimal value of $\beta$ decreases. This manifests in larger variance of the samples from the physics-informed prior, as evidenced in Figure 2, and can be interpreted as a lower level of trust in the assumed physics. On the other hand, if we have selected the perfect model, i.e. $\|\hat{u}^{*}-\hat{u}^{0}\|_{H_{k}}^{2}=0$ then $\beta\to\infty$. The prior then collapses to a Dirac centered at $u^{0}$, which signals the absence of model-form error.
@@
$\beta\sim p(\beta),\quad\hat{q}\sim p(\hat{q})$
$\hat{u}|\beta,\hat{q}\sim\mathcal{N}(\hat{u}^{0}(\cdot;\hat{q}),\beta^{-1}\Sigma_{\mathcal{F}})$
@@
$p(\beta,\hat{q}|d)\propto\mathcal{N}(d|\hat{u}^{0}(\cdot;\hat{q}),\beta^{-1}\Sigma_{\mathcal{F}}+\sigma^{2}I)p(\beta)p(\hat{q}).$ (5.6)
Observe how the variance of eq. (5.6) changes according to the estimate of $\beta$ in Theorem 5.1. A model with relatively high error will result in a smaller value of $\beta$. This can result from either choosing the wrong PDE, or by incorrectly identifying the source. In this situation, the variance in the prediction over $\hat{q}$ increases.
The intuition here is that if the model is wrong, the posterior obtained from the methodology responds with a lower confidence in the prediction of $\hat{q}$. Likewise, if the model-form error is low, the method becomes more confident about the prediction. This behavior is typically absent from Bayesian methods, as the posterior variance is invariant to model-form error. Also of note is that the posterior variance never entirely disappears due to the presence of measurement noise. This agrees with the usual result that identifying the source term of the Poisson equation is an ill-posed inverse problem [23].

### v2

@@ 4.4 A note on the use of neural networks
Our finite-dimensional representation was derived using a truncated orthonormal basis. It is natural to ask whether other parameterizations are also suitable, particularly for deep neural networks. For instance, this would provide a basis to connect our work to PINNs (more specifically deep Ritz). When working with neural networks, there are some technical issues which must be treated with care. We touch on both treating the the space $\mathcal{F}$ as a collection of neural networks as well as the convergence theorems.
Recall we look to approximate the prior with the finite-dimensional representation as given by eq. (4.2). This representation is not immediately well-defined if we parameterize $u$ with a neural network. To summarize, for a fixed neural network structure, we cannot assume that the space of functions the network can represent is finite-dimensional, in which case the Lebesgue measure would not exist. To explain this, we introduce some notation following [petersen2021topological, mahan2021nonclosedness].
Let $\Phi=\{(A_{\ell},b_{\ell})\}_{\ell=1}^{n_{L}}$ be a set of matrix-vector tuples where $A_{\ell}\in\mathbb{R}^{N_{\ell}\times N_{\ell-1}}$ and $b_{\ell}\in\mathbb{R}^{N_{\ell}}$ for each $\ell$. The architecture of the network is given by $S=(N_{0},N_{1},\dots,N_{n_{L}})$, where $N(S)$ is the total number of neurons and $n_{L}=n_{L}(S)$ is the number of layers. The collection $\Phi$ represents the values of the weights for a neural network with architecture $S$. Then, for an activation function $h:\mathbb{R}\to\mathbb{R}$, the neural network is given by a mapping $\mathrm{NN}_{h}(\Phi):\Omega\to\mathbb{R}$. We are interested in the properties of the function space induced by the network for fixed $S$ and $h$. We will denote this set by $\mathcal{R}(\mathrm{NN_{h}})(S)$.
As it turns out, if we allow $\Phi$ to vary arbitrarily, then $\mathcal{R}(\mathrm{NN_{h}})(S)$ is not closed in $L^{p}(\Omega)$, $0<p<\infty$, for all activation functions commonly used in PINNs [petersen2021topological, Theorem 3.1]. The same is true in Sobolev spaces [mahan2021nonclosedness]. This implies that the function space $\mathcal{F}=\mathcal{R}(\mathrm{NN_{h}})(S)$ is not finite-dimensional and eq. (4.2) is no longer well-defined.
However, if $\Phi$ is restricted to a compact set, then $\mathcal{R}(\mathrm{NN_{h}})(S)$ is compact in $L^{p}(\Omega)$ [petersen2021topological, Proposition 3.5]. This compact restriction of $\Phi$ results from schemes which prevent exploding weights, a common practice. While the result is nicer, it is still not immediately applicable to the construction of our finite-dimensional approximation: there is no guarantee that a compact set of a Hilbert space will be finite-dimensional. The Hilbert cube is one example. Although, we may approximate any compact set with a finite-dimensional subspace to arbitrary accuracy.
Theorem 4.8 ([stan2000paley]).
@@
Therefore, one could theoretically take $\mathcal{F}$ to be a finite-dimensional space which approximates $\mathcal{R}(\mathrm{NN}_{h})(S)$ to a given tolerance, $\varepsilon$. The size of $\mathcal{F}$ on which $\mu_{n}$ is defined may be adjusted by tweaking $\varepsilon$, changing the bound on the weights, or changing the network structure.
In Corollary 4.2, we show that the function which solves the physics-regularized inverse problem will converge to the ground truth in the large-data limit. This is if we solve the problem in the infinite-dimensional setting. Ideally we would like to derive the result for training neural networks. The solution to this problem will be a function which lives in the RKHS $H^{1}_{0}(\Omega)$. Again if we allow $\Phi$ to vary arbitrarily, then $\mathcal{R}(\mathrm{NN}_{h})(S)$ is not closed in $H^{1}_{0}(\Omega)$. This means that there are functions in $H^{1}_{0}(\Omega)$ for which the neural network must send $\|\Phi\|\to\infty$ in order to approximate. If the ground truth happens to be such a function, then the convergence theorem will not hold. Likewise, if we limit $\Phi$ to a compact set, then $\mathcal{R}(\mathrm{NN}_{h})(S)$ is compact in $H^{1}_{0}(\Omega)$. In this case, the neural network is only able to approximate a function to any accuracy if that function is also a neural network, so it is unlikely that convergence holds. The only case where convergence to the ground truth could hold is if we allow the architecture of the neural network to change arbitrarily so that we may rely on the universal approximation theorem [hornik1989multilayer]. However, establishing this connection is beyond the scope of our current work.
@@ 5 On model-form error
In this section we perform an analysis of the hyperparameter $\beta$ towards the application of detecting model-form error. Since $\beta$ is a hyperparameter of the GP prior, it is natural to assess how $\beta$ is learned during training. We show the optimal choice of $\beta$ adjusts according to model-misspecification. We build towards the result by working with the finite-dimensional distributions discussed in Section 4.2.
Start by introducing a finite-dimensional representation of $L^{2}(\Omega)$. This representation induces the function space $\mathcal{F}\subset L^{2}(\Omega)$ with $\dim(\mathcal{F})=M<\infty$. We will then study the posterior behavior of $\beta$ in the limit of infinite data. Given our training data of the form eq. (2.3), we begin by writing the problem down as a hierarchical model:
$\beta$ $\sim p(\beta),$
$\hat{u}|\beta$ $\sim\mathcal{N}(\hat{u}^{0},\beta^{-1}\Sigma_{\mathcal{F}_{M}}),$
@@
where $\hat{u}^{0}$ is the projection of $u^{0}$ onto $\mathcal{F}$, and $\Sigma_{\mathcal{F}_{M}}$ is the restriction of the covariance operator given by eq. (3.4) to $\mathcal{F}$. Since we are no longer in the infinite-dimensional setting, application of Bayes’s rule in the usual sense is justified, and we can also rely on the Lebesgue integral when deriving expressions. We derive the joint posterior
@@
$\beta^{*}=\operatorname*{arg\,max}_{\beta\in(0,\infty)}\log\int\frac{1}{Z}p(d|\hat{u})p(\hat{u}|\beta)d\hat{u}+\log p(\beta).$
Note that this is equivalent to maximizing the $\log$-evidence given by $\mathcal{L}(\beta)\coloneqq\log p(d|\beta)+\log p(\beta)$. In what follows, we will show the MAP estimate is unique in the limit of large data, for certain choices of $p(\beta)$. We start by deriving an expression for the gradient of the target function. Throughout, we will center the space so that the prior mean function becomes 0. We have shown this is valid as the prior mean function does not depend on $\beta$ and it also lives in $H_{k}$.
Lemma 5.1.
Consider the hierarchical model eq. (5.1). The $\log$-evidence of this model $\mathcal{L}(\beta)=\log p(d|\beta)+\log p(\beta)$ satisfies
$\frac{\partial}{\partial\beta}\mathcal{L}(\beta)=\frac{M}{2\beta}+\frac{\partial}{\partial\beta}\log p(\beta)-\frac{1}{2}\langle\tilde{m}-\hat{u}^{0},\Sigma^{-1}_{\mathcal{F}_{M}}(\tilde{m}-\hat{u}^{0})\rangle-\frac{1}{2}\mathrm{tr}\left(\Sigma^{-1}_{\mathcal{F}_{M}}\tilde{\Sigma}_{\mathcal{F}_{M}}\right),$
@@
The $\log\beta$ term is obvious. What remains is the marginal $\log$-likelihood term $\log p(d|\beta)$. We start by using Fisher’s identity, see [louis1982finding, eq. 3.1] or [cappe2005inference, Prop. 10.1.6], which relates the derivative of the marginal $\log$-likelihood to an expectation over the posterior. Letting $\partial_{\beta}$ denote the partial derivative with respect to $\beta$, we have
$\partial_{\beta}\log p(d|\beta)$ $=\mathbb{E}_{\hat{u}|d,\beta}\left[\partial_{\beta}\log p(\hat{u}|\beta)\right]$
$=\mathbb{E}_{\hat{u}|d,\beta}\left[\partial_{\beta}\left(\frac{M}{2}\log\beta-\frac{1}{2}\log\det\Sigma^{-1}_{\mathcal{F}_{M}}-\frac{\beta}{2}\langle\hat{u}-\hat{u}^{0},\Sigma^{-1}_{\mathcal{F}_{M}}(\hat{u}-\hat{u}^{0})\rangle\right)\right]$
$=\mathbb{E}_{\hat{u}|d,\beta}\left[\frac{M}{2\beta}-\frac{1}{2}\langle\hat{u}-\hat{u}^{0},\Sigma^{-1}_{\mathcal{F}_{M}}(\hat{u}-\hat{u}^{0})\rangle\right]$
$=\frac{M}{2\beta}-\frac{1}{2}\mathbb{E}_{\hat{u}|d,\beta}\left[\langle\hat{u}-\hat{u}^{0},\Sigma^{-1}_{\mathcal{F}_{M}}(\hat{u}-\hat{u}^{0})\rangle\right],$
which gives the first term. What remains is to compute the posterior expectation. Application of a standard Gaussian identity yields
$\mathbb{E}_{\hat{u}|d,\beta}\left[\langle\hat{u}-\hat{u}^{0},\Sigma^{-1}_{\mathcal{F}_{M}}(\hat{u}-\hat{u}^{0})\rangle\right]=\langle\tilde{m}-\hat{u}^{0},\Sigma^{-1}_{\mathcal{F}_{M}}(\tilde{m}-\hat{u}^{0})\rangle+\mathrm{tr}\left(\Sigma^{-1}_{\mathcal{F}_{M}}\tilde{\Sigma}_{\mathcal{F}_{M}}\right),$
which shows the remaining terms. ∎
Remark 5.1.
The expression we have derived in Lemma 5.1 for $\partial_{\beta}\mathcal{L}(\beta)$ can be seen elsewhere, but written in terms of a quadratic expression of the data $d$. For example, this appears in [williams2006gaussian, Equation 5.9] or in the function space setting, which is our case, in [dunlop2020hyperparameter] (although they do not derive the gradient). What we have done is state the result in a way in which the dependency of the $\log$-evidence on the posterior mean is made explicit. This enables us to asymptotically identify the unique estimate of $\beta$ through application of Theorem 4.6 and Theorem 4.7. Whereas the previous works do so only in numerical experiments.
We now identify the MAP estimate of $\beta$ in the large data limit under different prior choices.
@@
Let $\mathcal{L}(\beta)\coloneqq\log\int\frac{1}{Z}p(d|\hat{u})p(\hat{u}|\beta)d\hat{u}+\log p(\beta)$ as given by the hierarchical model eq. (5.1). Further, let the assumptions of Theorem 4.5 hold. Then, in the limit $h_{X_{n}}\to 0$, we have the following.
  1. 1.
    If $\beta$ is assigned a flat prior, then β^* = M^u*- ^u0, ΣFM-1(^u*-^u0).
  2. 2.
    If $\beta$ is assigned Jeffreys prior, then β^* = M -2 ^u*- ^u0, ΣFM-1(^u*-^u0).
@@
$\frac{\partial}{\partial\beta}\mathcal{L}(\beta)=\frac{M}{2\beta}+\frac{\partial}{\partial\beta}\log p(\beta)-\frac{1}{2}\langle\tilde{m}-\hat{u}^{0},\Sigma^{-1}_{\mathcal{F}_{M}}(\tilde{m}-\hat{u}^{0})\rangle-\frac{1}{2}\mathrm{tr}\left(\Sigma^{-1}_{\mathcal{F}_{M}}\tilde{\Sigma}_{\mathcal{F}_{M}}\right).$
Now, by Theorem 4.6 we have $\tilde{m}\to u^{*}$, and by Theorem 4.7 $\tilde{k}\to 0$ as $h_{X_{n}}\to 0$. Passing to the limit, the gradient becomes
$\frac{\partial}{\partial\beta}\mathcal{L}(\beta)=\frac{M}{2\beta}+\frac{\partial}{\partial\beta}\log p(\beta)-\frac{1}{2}\langle\hat{u}^{*}-\hat{u}^{0},\Sigma^{-1}_{\mathcal{F}_{M}}(\hat{u}^{*}-\hat{u}^{0})\rangle.$
@@
$\langle\hat{u}^{*}-\hat{u}^{0},\Sigma_{\mathcal{F}_{M}}^{-1}(\hat{u}^{*}-\hat{u}^{0})\rangle=\|\hat{u}^{*}-\hat{u}^{0}\|_{H_{k}}^{2},$
restricted to $\mathcal{F}$, which was derived in the proof of Proposition 3.2. That is, the optimal value of $\beta$ is inversely proportional to the RKHS distance between the prior mean function $u^{0}$ and the ground truth $u^{*}$. The same holds for $L^{2}(\Omega)$ distance by the Sobolev embedding theorem. Recall that $u^{0}$ is exactly the unique solution to the chosen physical model eq. (1.1). Hence, $\beta$ is sensitive to the distance between the true field $u^{*}$, and the one we have a priori assumed is correct $u^{0}$. As $u^{*}$ moves further away from $u^{0}$, the optimal value of $\beta$ decreases. This manifests in larger variance of the samples from the physics-informed prior, as evidenced in Fig. 1, and can be interpreted as a lower level of trust in the assumed physics. On the other hand, if we have selected the perfect model, i.e. $\|u^{*}-u^{0}\|_{H_{k}}^{2}=0$ then $\beta\to\infty$. The prior then collapses to a Dirac centered at $u^{0}$, which signals the absence of model-form error.
@@
[…]
$\hat{u}|\beta,\hat{q}\sim\mathcal{N}(\hat{u}^{0}(\cdot;\hat{q}),\beta^{-1}\Sigma_{\mathcal{F}_{M}}),$
@@
Observe how the variance of eq. (5.3) changes according to the MAP estimate of $\beta$ in Theorem 5.1. A model with relatively high error will result in a smaller value of $\beta$. This can result from either choosing the wrong PDE, or by incorrectly identifying the source. In this situation, the variance in the prediction over $\hat{q}$ increases.
The intuition here is that if the model is wrong, the posterior obtained from the methodology responds with a lower confidence in the prediction of $\hat{q}$. Likewise, if the model-form error is low, the posterior reflects a higher degree of confidence in the prediction. This behavior is typically absent from Bayesian methods, as the posterior variance is invariant to model-form error. Also of note is that the posterior variance never entirely disappears due to the presence of measurement noise. This agrees with the usual result that identifying the source term of the Poisson equation is an ill-posed inverse problem [ghattas2021learning].